# Quantum variational calculus on a lattice

### v1

@@ 3.1. Construction of the double complex $\Omega(J^{\infty})$
where we keep $i$ fixed and vary $f$ in the $\delta_{j}$ direction for the coefficient of $\mathrm{d}f_{j}$. Combining this with the above we can write this as
@@
[…]
@@
which indeed necessarily holds on applying $\mathrm{d}$ to $[u_{I},\phi]=0$. Similarly expanding
@@
[…]
@@
$C(J^{\infty})$ generated by functions on $\mathbb{Z}$ and $\{u_{I}\}$ as above extends to an exterior algebra $\Omega(J^{\infty})$ generated by $\Omega(\mathbb{Z})$ and additional generators $\mathrm{d}u_{I}$ with relations
@@
$[\mathrm{d}_{V}u_{I},\Phi]=0,\quad\{e^{a},\mathrm{d}_{V}u_{I}\}+\mathrm{d}_{V}u_{aI}\wedge e^{a}=0,\quad\{\mathrm{d}_{V}u_{I},\mathrm{d}_{V}u_{J}\}=0$
@@ 3.2. Euler-Lagrange equations
For $L=L(u,u_{a})$ we define (with sums over $a$),
@@
where in the second line we have used $-D_{a}=R_{a}D_{a^{-1}}$ and the property (12) for $R_{a}$. Writing now $R_{a}=D_{a}+\mathrm{id}$ results in
@@
In the continuum limit, where we approximate $\mathbb{Z}$ to $\mathbb{R}$ with coordinate $t$, we expect $D_{+}$ and $D_{-}$ to correspond to the positive and negative total derivatives in the $t$ direction ±$\pm D_{1}$, and hence we recover the classical EL form in Equation (1). Similarly we expect the shift operation $R_{a}$ to correspond to the identity in the limit, recovering the boundary form in (1). Note that the relation $D_{a}=R_{a}-\mathrm{id}$ now needs to be scaled so that the finite difference becomes a usual derivative, which then also enters $R_{a}u$.
As the bundle $\mathbb{Z}\times\mathbb{R}\to\mathbb{Z}$ models a point particle on $\mathbb{R}$ evolving in discrete time, we want to reproduce Example 2.1 in this setting, starting with a free particle $q=\phi:\mathbb{Z}\to\mathbb{R}$ with $V(q)=0$. Let $(e^{a},e^{b})=g^{ab}=\delta^{a,b^{-1}}$ represent the inverse Euclidean metric on $\mathbb{Z}$, and define the integral on $\mathbb{Z}$ as $\int_{\mathbb{Z}}f\mathrm{Vol}=\sum_{i}f_{i}$. Then in analogy to Example 2.1 we set the action to be
@@
$=\frac{m}{4}\int_{\mathbb{Z}}\left((\partial_{+}q)^{2}+(\partial_{-}q)^{2}\right)\mathrm{Vol}$
Where we have used $R_{a}\partial_{a^{-1}}=-\partial_{a}$. Note the prefactor $-\frac{m}{4}$ instead of ${m\over 2}$. The minus sign comes from the overall minus sign that appears in $(\mathrm{d}q,\mathrm{d}q)$, and the extra $1\over 2$ is introduced to ensure that the continuum limit is correct, since in the limit $\mathbb{Z}\to\mathbb{R}$ we expect ±$\partial_{\pm}q$ to correspond to ±$\pm\partial_{t}\phi$, and therefore $(\partial_{+}q)^{2}+(\partial_{-}q)^{2}$ to $2(\partial_{t}q)^{2}$. The Lagrangian is then $L={m\over 4}\sum_{a}u^{2}_{a}$, with EL and boundary forms given by
@@
In the continuum limit, we expect $\sum_{a}u_{aa^{-1}}$ and $u_{a^{-1}}$ to correspond to $-2u_{tt}$ and $-u_{t}$ respectively, thus recovering the results of Example 2.1. Using now $u_{aa^{-1}}=-u_{a}-u_{a^{-1}}$, we find the EL equation, or Newton’s equation, in discrete time to be
@@
We see that an unexpected extra term needs to be added to obtain a conserved current, namely $-{m\over 4}\epsilon^{b}u_{a^{-1}b}u_{b}\mathrm{Vol}_{a}$. The noncommutative-geometric origin of this term is not clear but may relate to our assumption that $\sigma=0$.
@@
where (,)(\ ,\ )( , ) is the metric inner product and in our case $\nabla$ obeys $\nabla e^{a}=0$ and the Leibniz rule and extends to tensor products in an obvious way (it is a bimodule connection in the sense used in [2] with the generalised braiding given by flip on the basis). Hence divergence-free just amounts to $\sum_{a}\partial_{a^{-1}}(T_{ab})=0$ for all $b$.
@@
$E[q]=-{m\over 2}(\partial_{+}q)\partial_{-}q$
The minus sign is important in order to recover the kinetic energy in the classical limit from Example 2.1. In the continuum limit we expect as discussed ±$\partial_{\pm}q$ to correspond to ±$\pm\partial_{t}q$, making that the sign of the above kinetic energy is positive as expected.
Since this is a new formalism, we check this claim explicitly. Using the modified Leibniz rule for finite differences,
$\partial_{+}((\partial_{+}q)(\partial_{-}q))$ =(+$=(\partial^{2}_{+}q)R_{+}(\partial_{-}q)+(\partial_{+}q)(\partial_{+}\partial_{-}q)$
@@
The current from Proposition 3.4 is conserved when a quadractic potential $V(q)=\mu q^{2}$ is included in the Lagrangian such that $L(u,u_{a})={m\over 4}\sum_{a}u^{2}_{a}-V(u)$.
@@
Repeating the analysis with the stress-energy tensor, the conserved charge now corresponds to the total energy
$E[q]=-{m\over 2}(\partial_{+}q)(\partial_{-}q)+V(q).$
It is not clear how to extend this result to higher order potentials.
@@
Next, for the Euler-Lagrange equations, we still have a unique (up to scale) top form ${\rm Vol}=e^{1}\wedge\cdots\wedge e^{|\mathcal{C}|}$ and we use this to define $\mathrm{Vol}_{a}$ as a product of the 1-forms without $e^{a}$ such that $\mathrm{Vol}=e^{a}\wedge\mathrm{Vol}_{a}$ (no sum). Then the computation of $\mathrm{d}_{V}(L\mathrm{Vol})$ and hence of the EL form and boundary form goes as in Theorem 3.2, wielding the same results, now with $a\in\mathcal{C}$. Similarly, Section 3.3 can be reproduced up to the introduction of a Lagrangian and correspondingly the form of the stress-energy tensor $T_{ab}$, but the computation of the conserved charges needs extra care, and is also a little different when we make the interpretation 1+0 classical field theory. We will give the results in detail for $\mathbb{Z}^{m},\mathbb{Z}^{1,m-1}$, but the method works in the same way for any Abelian group of the form ×$\mathbb{Z}\times G$ or $\mathbb{Z}_{N}\times G$, replacing $\mathcal{C}_{m-1}$ as defined below by $\mathcal{C}_{G}$.
@@ 4.1. EL equations and translation symmetries for $X=\mathbb{Z}^{m},\mathbb{Z}^{1,m-1}$
where the ±1$\pm$ 1± 1 is in the $i$-th position. The volume form is then
@@
where $e^{i\pm}$ corresponds to the direction ±$\pm v^{i}$.
@@
where we again introduce the factor $-{1\over 2}$ in front of the kinetic term in order to recover the classical action from Example 2.2. Recall that the minus cancels the one coming from $(\mathrm{d}\phi,\mathrm{d}\phi)$, and the ${1\over 2}$ is there since every classical direction corresponds to two directions in $\mathcal{C}$ (namely $e^{i\pm}$). Given the Lagrangian $L=\frac{1}{4}\sum_{a\in\mathcal{C}}u_{a}^{2}-\frac{1}{2}m^{2}u^{2}$ we can use Theorem 3.2 to compute the EL and boundary forms
@@
is the standard graph or lattice Laplacian with the Euclidean (constant) metric on edges and $\nabla(\mathrm{d}e_{a})=0$, cf[2]. We thus recover the Klein-Gordon equation $(\Delta+m^{2})\phi=0$ for this case. The 1/21/21 / 2 relates to the doubling of derivatives.
@@
$-\frac{1}{2}m^{2}u^{2}_{b}$ $=\frac{1}{2}D_{b}\left(\sum_{a\in\mathcal{C}}u_{aa^{-1}}-m^{2}u\right)u_{b}-\frac{1}{2}\sum_{a\in\mathcal{C}}u_{aa^{-1}b}u_{b}$
@@
$\mathrm{d}_{H}j_{0}=-\sum_{b\in\mathcal{C}}\epsilon^{b}\left(\left({1\over 2}\sum_{a\in\mathcal{C}}u_{aa^{-1}}-m^{2}u\right)u_{b}+\frac{1}{2}D_{b}\left({1\over 2}\sum_{a\in\mathcal{C}}u_{aa^{-1}}-m^{2}u\right)u_{b}-\frac{1}{4}\sum_{a\in\mathcal{C}}D_{a}(u_{a^{-1}b}u_{b})\right)\mathrm{Vol}.$
The first and second terms vanish when the EL equations are fulfilled, and the last can be written as $-\sum_{a,b\in\mathcal{C}}\frac{1}{2}\mathrm{d}_{H}(u_{a^{-1}b}u_{b})\mathrm{Vol}_{a})$, so that the current
@@
To see this we want to relate the divergence free condition for $b$ and $b^{-1}$. Comparing the expression for $\mathrm{d}_{H}j$ from Proposition 4.1 and $j=\sum_{a,b\in\mathcal{C}}\epsilon^{b}T_{ab}\mathrm{Vol}_{a^{-1}}$ we can relate the divergence of the stress-energy tensor to an expression which depends on the EL equations. From there we compute
@@
To consider conserved charges, we now split $\mathbb{Z}^{m}=\mathbb{Z}\times\mathbb{Z}^{m-1}$, with corresponding generating sets $\mathcal{C}_{1}=\{t,t^{-1}\}$ and $\mathcal{C}_{m-1}=\mathcal{C}\backslash\{\mathcal{C}_{1}\}$, where $t,t^{-1}$ are the positive and negative Euclidean time directions ±$\pm v^{1}$. The volume form on $\mathbb{Z}^{m-1}$ is then
@@
With this split we can define conserved charges in a similar way as in Corollary 2.4.
@@
is conserved in the sense that $\partial_{t}Q=0$ on-shell.
@@
vanishing on-shell. Using $\partial_{a}e_{\infty}^{*}(\Phi)=e_{\infty}^{*}(D_{a}\Phi)$, $-D_{a}R_{a^{-1}}=D_{a^{-1}}$ and $\int_{\mathbb{Z}^{m-1}}\mathrm{Vol}_{\mathbb{Z}^{m-1}}\partial_{b}=0$ for $b\in\mathcal{C}_{m-1}$, we have
@@
on-shell. ∎
@@
$E$ $=\int_{\mathbb{Z}^{m-1}}(T_{t^{-1}t}[\phi]-R_{t^{-1}}T_{tt}[\phi])\mathrm{Vol}_{\mathbb{Z}^{m-1}}$
$=-{1\over 2}\int_{\mathbb{Z}^{m-1}}\left((\partial_{t}\phi)(\partial_{t^{-1}}\phi)+{1\over 2}\sum_{b\in\mathcal{C}_{m-1}}(\partial_{b}\phi)^{2}-m^{2}\right)\mathrm{Vol}_{\mathbb{Z}^{m-1}},$
$P_{b}$ $=\int_{\mathbb{Z}^{m-1}}(T_{t^{-1}b}[\phi]-R_{t^{-1}}T_{tb}[\phi])\mathrm{Vol}_{\mathbb{Z}^{m-1}}$
$=-{1\over 2}\int_{\mathbb{Z}^{m-1}}(\partial_{t^{-1}}\phi)\left(\partial_{b}\phi-\partial_{b^{-1}}\phi\right)\mathrm{Vol}_{\mathbb{Z}^{m-1}}.$

### v2

@@ 3.1. Construction of the double complex $\Omega(J^{\infty})$
where we keep $i$ fixed and vary $\phi$ in the $\delta_{j}$ direction for the coefficient of $\mathrm{d}\phi^{j}$. Combining this with the above we can write this as
@@
where $aI$ denotes the standard lexicographic form of the indices with an extra $a$. For example, if $I=+$- then +I=+++I=++- and $I=+-$-. From this we conclude (summing $a$) that
@@
which indeed necessarily holds on applying $\mathrm{d}$ to $u_{I},\phi]=0$. Similarly expanding
@@
where $\{\ ,\ \}$ denotes anticommutator. This seems reasonable and is consistent with $\mathrm{d}_{H}=[\theta,\ \}$ being inner. This then implies that $\mathrm{d}_{V}^{2}=0$ on $u_{I}$. The $\mathrm{d}_{V},\mathrm{d}_{H}$ then extend by the graded-derivation rules to all $\Omega(\mathbb{Z}\times\mathbb{R}^{\infty})$ as one can check. Thus, we arrive at the following $\Omega(J^{\infty})$.
@@
$C(J^{\infty})$ generated by functions on $\mathbb{Z}$ and $\{u_{I}\}$ as above extends to an exterior algebra $\Omega(J^{\infty})$ generated by $\Omega(\mathbb{Z})$, the $u_{I}$ and additional generators $\mathrm{d}u_{I}$ with relations
@@
$\mathrm{d}_{V}u_{I},\Phi]=0,\quad\{e^{a},\mathrm{d}_{V}u_{I}\}+\mathrm{d}_{V}u_{aI}\wedge e^{a}=0,\quad\{\mathrm{d}_{V}u_{I},\mathrm{d}_{V}u_{J}\}=0$
@@ 3.2. Euler-Lagrange equations
For $L=L(u,u_{a})$, we define (with sums over $a$),
@@
where in the second line we have used $D_{a}=R_{a}D_{a^{-1}}$ and the property (13) for $R_{a}$. Writing now $R_{a}=D_{a}+\mathrm{id}$ results in
@@
In the continuum limit, where we approximate $\mathbb{Z}$ to $\mathbb{R}$ with coordinate $t$, we expect $D_{+}$ and $D_{-}$ to correspond to the positive and negative total derivatives in the $t$ direction $\pm D_{1}$, and hence we recover the classical EL form in Equation (2). Similarly we expect the shift operation $R_{a}$ to correspond to the identity in the limit, recovering the boundary form in (2). Note that the relation $D_{a}=R_{a}-\mathrm{id}$ now needs to be scaled so that the finite difference becomes a usual derivative, which then also enters $R_{a}u$.
As the bundle $\mathbb{Z}\times\mathbb{R}\to\mathbb{Z}$ models a point particle moving in $\mathbb{R}$ evolving in discrete time, we want to reproduce Example 2.1 in this setting, starting with a free particle $q=\phi:\mathbb{Z}\to\mathbb{R}$ with $V(q)=0$. Let $(e^{a},e^{b})=g^{ab}=\delta^{a,b^{-1}}$ represent the inverse Euclidean metric on $\mathbb{Z}$, and define the integral on $\mathbb{Z}$ as $\int_{\mathbb{Z}}f\mathrm{Vol}=\sum_{i}f_{i}$. Then in analogy to Example 2.1 we set the action to be
@@
$=\frac{m}{4}\int_{\mathbb{Z}}\left((\partial_{+}q)^{2}+(\partial_{-}q)^{2}\right)\mathrm{Vol},$
where we have used $R_{a}\partial_{a^{-1}}=-\partial_{a}$. Note the prefactor $\frac{m}{4}$ instead of $m\over 2$}. The minus sign comes from the overall minus sign that appears in $(\mathrm{d}q,\mathrm{d}q)$, and the extra $1\over 2$ is introduced to ensure that the continuum limit is correct, since in the limit $\mathbb{Z}\to\mathbb{R}$ we expect $\partial_{\pm}q$ to correspond to $\pm\partial_{t}\phi$, and therefore $(\partial_{+}q)^{2}+(\partial_{-}q)^{2}$ to $2(\partial_{t}q)^{2}$. The Lagrangian is then $L={m\over 4}\sum_{a}u^{2}_{a}$, with EL and boundary forms given by
@@
In the continuum limit, we expect $\sum_{a}u_{aa^{-1}}$ and $u_{a^{-1}}$ to correspond to $2u_{tt}$ and $u_{t}$ respectively, thus recovering the results of Example 2.1. Using now $u_{aa^{-1}}=-u_{a}-u_{a^{-1}}$, we find the EL equation, or Newton’s equation, in discrete time to be
@@
We see that an unexpected extra term needs to be added to obtain a conserved current, namely ${m\over 4}\epsilon^{b}u_{a^{-1}b}u_{b}\mathrm{Vol}_{a}$. The noncommutative-geometric origin of this term is not clear but may relate to our assumption that $\sigma=0$.
@@
where (,)(\ ,\ ) is the metric inner product and in our case $\nabla$ obeys $\nabla e^{a}=0$ and the Leibniz rule and extends to tensor products in an obvious way (it is a bimodule connection in the sense used in [2] with the generalised braiding given by flip on the basis). Hence divergence-free just amounts to $\sum_{a}\partial_{a^{-1}}(T_{ab})=0$ for all $b$.
@@
$E[q]=-{m\over 2}(\partial_{+}q)(\partial_{-}q)$
The minus sign is important in order to recover the kinetic energy in the classical limit from Example 2.1. In the continuum limit we expect as discussed $\partial_{\pm}q$ to correspond to $\pm\partial_{t}q$, making that the sign of the above kinetic energy is positive as expected.
Since this is a new formalism, we check the on-shell conservation claim explicitly. Using the modified Leibniz rule for finite differences,
$\partial_{+}((\partial_{+}q)(\partial_{-}q))$ $=(\partial^{2}_{+}q)R_{+}(\partial_{-}q)+(\partial_{+}q)(\partial_{+}\partial_{-}q)$
@@
The current from Proposition 3.4 is conserved when a quadractic potential $V(q)=\mu q^{2}$ is included in the Lagrangian $L(u,u_{a})={m\over 4}\sum_{a}u^{2}_{a}-V(u)$.
@@
Figure 1. Discrete time harmonic oscillator for $m=1$ and energy $E=1$ with increasing $\omega$. For the sine modes with $\omega<2$, the dashed line is the amplitude $q_{2}$ such that $E={1\over 8}m\omega^{2}(4-\omega^{2})q_{2}^{2}$, with minimum amplitude for $\omega=\sqrt{2}$. The period $2\pi/\omega^{\prime}$ is significantly smaller than $2\pi/\omega$ as $\omega$ increases. As $\omega\to 2$ from below, the amplitude blows up and the period tends towards the minimum alternating case. We also see a beat frequency emerging. The $\omega>2$ case has alternating exponential growth. Plots have been smoothly interpolated for visualisation.
Repeating the analysis with the stress-energy tensor, the charge now corresponds to the total energy
(16) $E[q]=-{m\over 2}(\partial_{+}q)(\partial_{-}q)+V(q)$
conserved on-shell for the equations of motion
(17) $\Delta_{\mathbb{Z}}q=-{1\over m}{\partial V(q)\over\partial q}$
when the potential is quadratic. It is not clear how to extend these results to higher order potentials, but we are able to construct a discrete time harmonic oscillator and we now look at this in more detail.
First of all, the left hand side of (17) is the standard lattice double-differential and writing $V(q)={1\over 2}{m\omega^{2}}q^{2}$ for frequency $\omega$, the equation becomes $\Delta_{\mathbb{Z}}q=-\omega^{2}q$. The solution to this on a lattice is well-known to be
$q(i)=q_{1}\cos(\omega^{\prime}i)+q_{2}\sin(\omega^{\prime}i);\quad\omega^{\prime}=\arccos\left(1-{\omega^{2}\over 2}\right)=\omega+{\omega^{3}\over 24}+O(\omega^{5})$
provided $\omega<2$. A short calculation gives energy
(18) $E[q]=\frac{m}{2}\sin^{2}(\omega^{\prime})(q_{1}^{2}+q_{2}^{2})={m\over 8}\omega^{2}(4-\omega^{2})(q_{1}^{2}+q_{2}^{2})$
which we see is $\geq 0$ and zero only for the zero solution. The mix of sine and cosine modes depends on the parameters and here is instructive to take them to be the values of $q(0)$ and $E(q)$. One can analyse this analytically or just go back to the equations of motion, solved by recursion with
$q(1)={q(0)\over 2}\left(2-\omega^{2}+\sqrt{8{E(q)\over mq(0)^{2}}-4\omega^{2}+\omega^{4}}\right)$
From this we see that there are real solutions if and only if
(19) $E[q]\geq E_{c}:={m\over 8}\omega^{2}(4-\omega^{2})q(0)^{2}.$
This is in line with classical mechanics where the total energy $E$ cannot be less than the potential energy at any point (here at $i=0$). For example, we can take $q(0)=0$ then $E_{c}=0$ and we see only sine solutions increasing in amplitude with increasing $E>0$ according to $E={1\over 8}m\omega^{2}(4-\omega^{2})q_{2}^{2}$ from (18). This modifies the relation $E={1\over 2}m\omega^{2}q_{2}^{2}$ in classical mechanics, parallel to the familiar fact in lattice field theory that the dispersion relation gets modified when $\omega$ is not small. Also parallel to the field theory case, if we let $\omega^{2}$ be negative then there are (real) unbounded exponential solutions. As we increase $\omega$, the amplitude for a given energy decreases until a minimum at $\omega=\sqrt{2}$, while the period $2\pi/\omega^{\prime}$ shrinks towards period 2, i.e alternating and blowing up in amplitude as $\omega\to 2$ from below. The case $\omega>2$ is not physical (if the discretisation is a model of Planck scale effects then these are transplankian modes) and here solutions blow up exponentially while alternating in $i$ (the ‘maximum frequency’). These effects are illustrated in Figure 1. Note that $E<E_{c}$ leads to complex oscillatory solutions but these are not part of our theory, and indeed the conserved energy for the complex case (which would be relevant to the scalar field theory point of view) is different as it involves conjugation of one of the fields.
Finally, the stress energy tensor, e.g. for the sine solutions $q(i)=q_{2}\sin(i\omega^{\prime})$ is
$T_{++}[q]$ $=-\frac{m}{4}q_{2}^{2}(\sin(i\omega^{\prime})-\sin((i+1)\omega^{\prime}))(\sin(i\omega^{\prime})-\sin((i+2)\omega^{\prime})),$
$T_{+-}[q]$ $=-\frac{m}{8}q_{2}^{2}(-2\sin(\omega^{\prime})\sin(2i\omega^{\prime})+\cos(2i\omega^{\prime})-\cos(2(i+1)\omega^{\prime})+\cos(2\omega^{\prime})-1),$
$T_{-+}[q]$ $=-\frac{m}{8}q_{2}^{2}\left(2\sin(\omega^{\prime})\sin(2i\omega^{\prime})-2\cos^{2}((i-1)\omega^{\prime})+\cos(2i\omega^{\prime})+\cos(2\omega^{\prime})\right),$
$T_{--}[q]$ $=\frac{m}{4}q_{2}^{2}(\sin((i-2)\omega^{\prime})-\sin(i\omega^{\prime}))(\sin(i\omega^{\prime})+\sin(\omega^{\prime}-i\omega^{\prime})),$
and is not itself constant in $i$ (but is divergence free by our construction).
@@
Next, for the Euler-Lagrange equations, we still have a unique (up to scale) top form ${\rm Vol}=e^{1}\wedge\cdots\wedge e^{|\mathcal{C}|}$ and we use this to define $\mathrm{Vol}_{a}$ as a product of the 1-forms without $e^{a}$ such that $\mathrm{Vol}=e^{a}\wedge\mathrm{Vol}_{a}$ (no sum). Then the computation of $\mathrm{d}_{V}(L\mathrm{Vol})$ and hence of the EL form and boundary form goes as in Theorem 3.2, wielding the same results, now with $a\in\mathcal{C}$. Similarly, Section 3.3 can be reproduced up to the introduction of a Lagrangian and correspondingly the form of the stress-energy tensor $T_{ab}$, but the computation of the conserved charges needs extra care, and is also a little different when we make the interpretation 1+0 classical field theory. We will give the results in detail for $\mathbb{Z}^{m},\mathbb{Z}^{1,m-1}$, but the method works in the same way for any Abelian group of the form $\mathbb{Z}\times G$ or $\mathbb{Z}_{N}\times G$, replacing $\mathcal{C}_{m-1}$ as defined below by $\mathcal{C}_{G}$.
@@ 4.1. EL equations and translation symmetries for $X=\mathbb{Z}^{m},\mathbb{Z}^{1,m-1}$
where the $\pm 1$ is in the $i$-th position. The volume form is then
@@
where $e^{i\pm}$ corresponds to the direction $\pm v^{i}$.
@@
where we again introduce the factor $1\over 2$} in front of the kinetic term in order to recover the classical action from Example 2.2. Recall that the minus cancels the one coming from $(\mathrm{d}\phi,\mathrm{d}\phi)$, and the $1\over 2$} is there since every classical direction corresponds to two directions in $\mathcal{C}$ (namely $e^{i\pm}$). Given the Lagrangian $L=\frac{1}{4}\sum_{a\in\mathcal{C}}u_{a}^{2}-\frac{1}{2}m^{2}u^{2}$ we can use Theorem 3.2 to compute the EL and boundary forms
@@
is the standard graph or lattice Laplacian with the Euclidean (constant) metric on edges and $\nabla(\mathrm{d}e_{a})=0$, cf. [2]. We thus recover the Klein-Gordon equation $(\Delta+m^{2})\phi=0$ for this case. The $1/2$ relates to the doubling of derivatives.
@@
$-\frac{1}{2}m^{2}u^{2}_{b}$ $=\frac{1}{2}D_{b}\left(\frac{1}{2}\sum_{a\in\mathcal{C}}u_{aa^{-1}}-m^{2}u\right)u_{b}-\frac{1}{4}\sum_{a\in\mathcal{C}}u_{aa^{-1}b}u_{b}$
@@
$\mathrm{d}_{H}j_{0}$ $=-\sum_{b\in\mathcal{C}}\epsilon^{b}\Big(\big({1\over 2}\sum_{a\in\mathcal{C}}u_{aa^{-1}}-m^{2}u\big)u_{b}+\frac{1}{2}D_{b}\big({1\over 2}\sum_{a\in\mathcal{C}}u_{aa^{-1}}-m^{2}u\big)u_{b}$
$\qquad\qquad\quad-\frac{1}{4}\sum_{a\in\mathcal{C}}D_{a}(u_{a^{-1}b}u_{b})\Big)\mathrm{Vol}.$
The first and second terms vanish when the EL equations are fulfilled, and the last can be written as $\sum_{a,b\in\mathcal{C}}\frac{1}{2}\mathrm{d}_{H}(u_{a^{-1}b}u_{b})\mathrm{Vol}_{a})$, so that the current
@@
To see this, we relate the divergence free conditions for $b$ and $b^{-1}$. Comparing the expression for $\mathrm{d}_{H}j$ from Proposition 4.1 and $j=\sum_{a,b\in\mathcal{C}}\epsilon^{b}T_{ab}\mathrm{Vol}_{a^{-1}}$ we can relate the divergence of the stress-energy tensor to an expression which depends on the EL equations. From there we compute
@@
[…]
@@
With this split, we can define conserved charges in a similar way to Corollary 2.4.
@@
is conserved in the sense that $\partial_{t}Q=0$ on-shell. Furthermore, the charge and current densities
$\rho[\phi]\coloneqq e_{\infty}^{*}(j^{t}-R_{t^{-1}}j^{t^{-1}})(\cdot,\phi),\quad\quad J^{x}[\phi]\coloneqq e_{\infty}^{*}(j^{x}-R_{x^{-1}}j^{x^{-1}})(\cdot,\phi)$
for $x$ the positive spatial directions in $\mathbb{Z}^{m-1}$, satisfy the continuity equation
$\partial_{t}\rho+\sum_{x}\partial_{x}J^{x}=0$
on-shell.
@@
vanishing on-shell. Using $\partial_{a}e_{\infty}^{*}(\Phi)=e_{\infty}^{*}(D_{a}\Phi)$, $D_{a}R_{a^{-1}}=D_{a^{-1}}$ and $\int_{\mathbb{Z}^{m-1}}\mathrm{Vol}_{\mathbb{Z}^{m-1}}\partial_{b}=0$ for $b\in\mathcal{C}_{m-1}$, we have
@@
on-shell. The continuity equation can be derived by rewriting the conservation of $j$ as
$D_{t}(j^{t}-R_{t^{-1}}j^{t^{-1}})+\sum_{x\in\mathbb{Z}_{m-1}}D_{x}(j^{x}-R_{x^{-1}}j^{x^{-1}}).$
@@
$E[\phi]$ $=\int_{\mathbb{Z}^{m-1}}(T_{t^{-1}t}[\phi]-R_{t^{-1}}T_{tt}[\phi])\mathrm{Vol}_{\mathbb{Z}^{m-1}}$
$=\int_{\mathbb{Z}^{m-1}}{1\over 2}\left(-(\partial_{t}\phi)(\partial_{t^{-1}}\phi)-{1\over 2}\sum_{b\in\mathcal{C}_{m-1}}(\partial_{b}\phi)^{2}+m^{2}\phi^{2}\right)\mathrm{Vol}_{\mathbb{Z}^{m-1}},$
$P_{b}[\phi]$ $=\int_{\mathbb{Z}^{m-1}}(T_{t^{-1}b}[\phi]-R_{t^{-1}}T_{tb}[\phi])\mathrm{Vol}_{\mathbb{Z}^{m-1}}$
$=\int_{\mathbb{Z}^{m-1}}{1\over 2}(\partial_{t^{-1}}\phi)\left(\partial_{b^{-1}}\phi-\partial_{b}\phi\right)\mathrm{Vol}_{\mathbb{Z}^{m-1}}.$